OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 5x - 3y - z if x = -2, y = 2, and z = -3

OpenStudy (anonymous):

You will substitute the given value in the variables. 5 (-2)-3(2)-(-3)=-10-6+3=-16+3=-13